OpenStudy (anonymous):

The integers differ by 12 and the sum of their squares is 74.

OpenStudy (lgbasallote):

let x and y be the integers x - y = 12 x^2 + y^2 = 74 solve for x and y does that help?

OpenStudy (anonymous):

x^2+y^2=74 then (y+12)^2+y^2=74 but how do i dothe next step?

OpenStudy (lgbasallote):

expand (y+12)^2

OpenStudy (anonymous):

use (a+b)^2=a^2+b^+2ab where a=y and b=12 now put into equ.

OpenStudy (anonymous):

use (a+b)^2=a^2+b^2+2ab where a=y and b=12 now put into equ.